Circular No. 6529

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996cc IN NGC 5673 P. Garnavich and R. Kirshner, Harvard-Smithsonian Center for As-

trophysics, write: “Spectra obtained by P. Berlind on Jan. 2.5 UT with the 1.5-m Tillinghast telescope show this object to be a type II supernova. Strong, broad, Hα emission is present with a shallow P-Cyg absorption trough. A narrow emission line of Hα from the host galaxy has a redshift of 2050 km/s.”

SUPERNOVA 1996cb IN NGC 3510 Garnavich and Kirshner report that Berlind also obtained spectra of

this object on Jan. 2.4 UT: “These show that an absorption feature has developed at 651.5 nm. The feature is likely to be He I 667.8-nm with an expansion velocity of 8000 km/s. He I 706.5-nm and 587.5-nm are also present with velocities near 8000 km/s. No He I features were present on a spectrum taken 1996 Dec. 17. The 1997 Jan. 2 spectrum is similar to that of the ‘IIb’ supernova 1993J in late-April 1993 (Filippenko et al. 1993, Ap.J. 415, L106). However, the Hα absorption in SN 1996cb is currently stronger than He I 587.5 nm. SN 1996cb may have lost a substantial fraction of its hydrogen envelope and evolve in a similar manner to SN 1993J. Radio detection by Van Dyk et al. (IAUC 6528) may indicate an early circumstellar interaction, supporting the classification as type ‘IIb’.”

R. Evans, Coonabarabran, N.S.W., who picked up this supernova at mag ∼ 14 on Jan. 3.7 UT during his visual supernova patrol, notes that it was not present when he last inspected NGC 3510 on 1996 Dec. 10.

NOTICE TO CONTRIBUTORS Effective immediately, there will be an increase in the line-charge rate

for those items published on these Circulars that are subject to these charges. The new charge will be $20.00 per line (in the printed version) plus $50.00 per item (the latter reduced to $40.00 for shared items). The items subject to line charges are generally those relating to other than discovery and initial confirmation of comets, novae and supernovae. Continuing astrometric and orbital data on comets and initial spectroscopic data on novae and supernovae are also exempt, as are contributions (not all of which are accepted for publication) by amateur astronomers.

1997 January 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden